OpenStudy (anonymous):

The radius of a circular puddle is growing at a rate of 5 cm/s. (a) How fast is its area growing at the instant when the radius is 10 cm? (b) How fast is the area growing at the instant when it equals 25 cm2?

zepdrix (zepdrix):

\[\large r'=5 \;cm/s\]Area of a circle is,\[\large A=\pi r^2\]Part a) is asking,\[\large \text{When r=10 cm find A'}\]

zepdrix (zepdrix):

Understand where the r' came from? :O That's the RATE OF CHANGE of r, that they gave us at the start.

zepdrix (zepdrix):

Are you more comfortable doing these problem written in Leibniz notation? Whatever works for you :D\[\large \frac{dr}{dt}=5 \; cm/s\]

OpenStudy (anonymous):

no i understand the other way. so i would just have to calculate to find A. and how would i find the answer to B?

zepdrix (zepdrix):

Yah for A we just take the derivative of the area function, take the derivative with respect to TIME, so don't forget that an r' will pop out when you apply the chain rule on the r term. Understand that part or should we go through it? :D

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

actually can you go through on how to find the derivative after calculating A= pie(10)^2

zepdrix (zepdrix):

We want to make sure we have BOTH of our VARIABLES in the problem when we differentiate. Don't plug anything in before differentiating! :O \[\large A=\pi r^2\]Taking the derivative WRT time gives us,\[\large A=\pi(2r)r'\]

zepdrix (zepdrix):

A' should be at the start of that bottom line :3 my bad.

zepdrix (zepdrix):

\[\large A'=\pi (2r)r'\]

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

now i would plug in 10 or 5 for r?

zepdrix (zepdrix):

You plug in these values, and solve for A'. \[\large r=10, \quad r'=5\]

OpenStudy (anonymous):

100(pie)

OpenStudy (anonymous):

which would be 314 after calculations

zepdrix (zepdrix):

Mmmmmmm yah looks good.

OpenStudy (anonymous):

im assuming its the same for b?

OpenStudy (anonymous):

actually b wants to find out the growing rate when it EQUALS 25cm

zepdrix (zepdrix):

Part B is a little tricky, right now, we have our Area function written in terms of r. r is the INDEPENDENT variable. We can plug in any value for the Radius, and from there we can calculate an Area. In part B, we are given a specific Area, so we're treating the area as the independent variable. So let's rewrite our function in terms of A.

zepdrix (zepdrix):

We want it to look like, r=something.

zepdrix (zepdrix):

Err actually... let's just find a radius that corresponds to A=25 cm. Then do it the normal way, that will probably be easier :)

zepdrix (zepdrix):

Plug in A=25, then solve for r.

zepdrix (zepdrix):

\[\large A=\pi r^2 \quad \rightarrow \quad 25=\pi r^2\] \[\large \frac{25}{\pi}=r^2 \quad \rightarrow \quad r=\sqrt{\frac{25}{\pi}} \quad \rightarrow \quad r=\frac{5}{\sqrt {\pi}}\]Something like that

zepdrix (zepdrix):

So now you have the R that corresponds to A=25. And now you can proceed the same why you did in part A, just with a new R value this time.

OpenStudy (anonymous):

ok how would i calculate 2(5/squareroot(pie)?

zepdrix (zepdrix):

What do you mean? Like how on a calculator? :O

zepdrix (zepdrix):

Oh, just multiply the 2 times the 5.

zepdrix (zepdrix):

\[\large A'=\pi(2r)r', \qquad r=\frac{5}{\sqrt{\pi}},\qquad r'=5\]\[\large A'=\pi\left(2\frac{5}{\sqrt{\pi}}\right)5 \quad \rightarrow \quad A'=\frac{50 \pi}{\sqrt{\pi}}\] You should get something like this when you plug everything in.. I think :O Where you getting stuck?

OpenStudy (anonymous):

thats what i got, but not sure if thats the final answer.